Lawrence Harris Department of Mathematics University of Kentucky Lexington, Kentucky 40506-0027 USA. E-mail address: larry@ms.uky.edu

Holomorphic Functions on Domains in Operator Spaces

Abstract: This talk will give an overview of facts and unsolved problems about bounded symmetric domains, Siegel domains of genus 2 and unbounded homogeneous domains in spaces of operators.